A $3500 loan was settled ten years later with a payment of $8100. Determine the annual rate of return.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

8100=3500*(1+r/100)^10

(8100/3500)^(1/10)=(1+r/100)

(1+r/100)=1.0875

r=1.0875-1

=8.75%(Approx).